0 0
Add a comment Improve this question Transcribed image text
Answer #1
the forces acting on the wall from both the sides are:
1/2*ρ1gh1*h1 from the left side where h1 = 7m (given) ata position h1/3 from the base.
1/2*ρ2gh2*h2 from the right side where h2 = h at aposition h2/3 from the base.
where ρ1 and ρ2 are the densities of seawater and fresh water
Note : we have not considered the force caused atmospheric pressurebecause, it will cancel out from both the sides.
so, to get a force equilibrium,
both these force should cancel each other
so, 1/2*ρ1*g*h12 =1/2*ρ2*g*h22
or, h2 = h1√(ρ1/ρ2)
or, h = 7√(ρ1/ρ2)
net moment M = 1/2*ρ1*g*h12*h1/3 -1/2*ρ2*g*h22*h2/3      in clockwisedirection
= g/6(ρ1*h13 - ρ2*h23)
= g/6(ρ1*73 -ρ2*73*ρ1/ρ2*√(ρ1/ρ2))
=(g/6)*ρ1*73(1-√(ρ1/ρ2))
we can put the values of ρ1= 1025 kg/m3 and ρ2 =1000 kg/m3
h = 7*√(1025/1000) = 7.08 m
M = (9.81/6)*1025*73*(1-√(1025/1000)) =-7140.95 N-m
i.e. 7140.95 N-m in counterclockwise direction.
Answer #2

Draw a long vertical wall separating the sea water and fresh water YF Freshwater Seawater Im Consider that the resultant forc